The angle between two vectors A and B each of magnitude x is 60 degree the values find the values of the scalar constant C such that a + c b has a magnitude of 2.18 x

Grade:11

1 Answers

Vikas TU
7 years ago
Let a = x1 i + x2 j + x3 k (cap)
       cb = c(y1 i + y2 j + y3 k) (cap)
 of a + cb = root((x1 +cy1)^2 + (x2 + cy2)^2 + (x3 + cy3)^2) = 2.18xmagnitude 
sq. both sides,
 
Also given, x^2 = root(x1^2 +x2^2 + x3^2) = root(y1^2 + y2^2 + y3^2)
also given
cos 60 = a.b/x^2 = 0.5
 
using all these relations solve the eqns. and find the a and b vector.